OpenStudy (anonymous):

I'll give you a medal and be a fan if you HELP PLEASE 1. The pump used to fill the water tank on a fire truck takes 40 minutes to fill the tank. If the pump sending water to the hose is on, it takes 60 minutes to fill the tank. If the pump filling the water tank is off, how long does the pump sending water to the hose take to empty the tank? A. 120 minutes B. 60 minutes C. 30 minutes D. 24 minutes 2. Annmarie can plow a field in 240 minutes. Gladys can plow a field 80 minutes faster. If they work together, how many minutes does it take them to plow the field? A. 96 minutes B. 160 minutes C. 400 minutes D. 480 minutes 3.Brody can fill a bowl with candy in 3 minutes. While Brody fills the bowl, Hudson takes the candy out of the bowl. With Hudson taking candy out of the bowl, it takes 5 minutes for Brody to fill the bowl. Which of the following can be used to determine the amount of time it takes for Hudson to empty the bowl if Perry does not add candy? A. 1 over 3 minus 1 over x equals 1 over 5 B. 1 over 3 minus 1 over 5 equals 1 over x C. 1 over 5 minus 1 over x equals 1 over 3 D. 1 over 3 minus 1 over x equals x over 5 4. Leanne can clean a fish tank in 30 minutes. Karl can clean a fish tank in 20 minutes. If they work together, how many minutes does it take them to clean a fish tank? A. 6 minutes B.10 minutes C. 12 minutes D. 50 minutes 5.Genna can build a cabinet in 4 hours. Claude can build a cabinet in only 3 hours. Which of the following can be used to determine the amount of time it would take for Genna and Claude to build a cabinet together? A. 1 over 4 plus 1 over x equals 1 over 3 B. 1 over 4 plus 1 over 3 equals 1 over x C. 1 over x plus 1 over 3 equals 1 over 4 D. 1 over 3 plus 1 over 4 equals x over 7

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (michele_laino):

Question 5) the working rate of Genna is W/4, the working rate of Claude is W/3, where W is the work to do. When Gwenna and Claude work together, then the working rate is: W/4 + W/3 so we can write: \[\Large \left( {\frac{W}{4} + \frac{W}{3}} \right)t = W\] where t is time needed to Genna and Claude to do the work W Please solve that equation for T, what do you get?

OpenStudy (michele_laino):

oops.. for t, what do you get?

OpenStudy (anonymous):

3 hrs?

OpenStudy (michele_laino):

hint: what is: \[\frac{1}{4} + \frac{1}{3} = ...?\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 12 }\]

OpenStudy (michele_laino):

ok! so, from preceding equation, we can write: \[\frac{{7W}}{{12}}t = W\] or: \[\frac{7}{{12}}t = 1\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

namely: \[t = \frac{{12}}{7}\]

OpenStudy (anonymous):

so the answer is D then?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thanks ! can you help with the others ?

OpenStudy (michele_laino):

ok! Question 4)

OpenStudy (michele_laino):

the working rate of Leanne is W/30, where the working rate of Karl is W/20, where as usual, W is the work to do. So when Leanne and Karl work together the working rate is: W/30+ W/20, and we can write: \[\Large \left( {\frac{W}{{30}} + \frac{W}{{20}}} \right)t = W\] where t is time neeeded to Leanne and Karl to finish the work W. Please solve for t

OpenStudy (michele_laino):

oops...whereas the working rate of Karl

OpenStudy (michele_laino):

as before, we can write: \[\Large \left( {\frac{1}{{30}} + \frac{1}{{20}}} \right)t = 1\]

OpenStudy (michele_laino):

what is t?

OpenStudy (anonymous):

\[\frac{ 5 }{ 15 } ?\]

OpenStudy (michele_laino):

we have: \[\frac{1}{{30}} + \frac{1}{{20}} = \frac{{2 + 3}}{{60}} = ...?\]

OpenStudy (anonymous):

i entered the wrong thing my bad!

OpenStudy (anonymous):

\[\frac{ 1 }{ 12 }\]

OpenStudy (michele_laino):

ok! So we have: \[\frac{t}{{12}} = 1\]

OpenStudy (michele_laino):

and therefore: \[t = 12 \times 1 = ...?\]

OpenStudy (anonymous):

12?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

it is 12 minutes

OpenStudy (anonymous):

Ok Thanks!

OpenStudy (anonymous):

Have time for another one?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

question 3)

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

I'm pondering....

OpenStudy (michele_laino):

the working rate of Brody is: W/3, when he works alone whereas the working rate of Brody is W/5, when Brody and Hudson work together W is the work to be done Now the working rate of Hudson is: W/x, where x is the unknown quantity. So, we can write: \[\Large \frac{W}{3} - \frac{W}{x} = \frac{W}{5}\] or: \[\Large \frac{1}{3} - \frac{1}{x} = \frac{1}{5}\] so, what is the right option?

OpenStudy (michele_laino):

hint: \[\Large \frac{1}{x} = \frac{1}{3} - \frac{1}{5}\]

OpenStudy (michele_laino):

@amielli

OpenStudy (anonymous):

2/15

OpenStudy (michele_laino):

yes! What is the right option?

OpenStudy (anonymous):

D?

OpenStudy (michele_laino):

I think that option D is wrong

OpenStudy (michele_laino):

please look at this equation: \[\Large \frac{1}{3} - \frac{1}{x} = \frac{1}{5}\] do you recognize to which option corresponds to?

OpenStudy (anonymous):

yes that would be A

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

I dont get it though

OpenStudy (michele_laino):

since Brody is filling the bowl, he is adding candies to that bowl. Hudson is subtracting candies to that bowl. Here is why I used the minus sign in the subsequent equation: \[\large \frac{W}{3} - \frac{W}{x} = \frac{W}{5}\]

OpenStudy (michele_laino):

now, is it clear?

OpenStudy (anonymous):

Oh yes I do

OpenStudy (anonymous):

Thanks

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

question 2)

OpenStudy (michele_laino):

here we can write: working rate of Annmarie is W/240, whereas the working rate of Gladies is W/80. Now when Annmarie and Gladies work together, the working rate is: W/240 + W/80. As usual W is the work to be done. So we can write: \[\Large \left( {\frac{W}{{240}} + \frac{W}{{80}}} \right)t = W\] or: \[\Large \left( {\frac{1}{{240}} + \frac{1}{{80}}} \right)t = 1\] Please solve for t

OpenStudy (michele_laino):

hint: \[\Large \frac{1}{{240}} + \frac{1}{{80}} = \frac{{1 + 3}}{{240}} = ...?\]

OpenStudy (michele_laino):

please wait a moment, I think to have made an error

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

the working rate of Gladies is: \[\Large \frac{W}{{240 - 80}} = \frac{W}{{160}}\]

OpenStudy (michele_laino):

so the right equation is: \[\Large \left( {\frac{W}{{240}} + \frac{W}{{240 - 80}}} \right)t = W\]

OpenStudy (michele_laino):

or: \[\Large \left( {\frac{1}{{240}} + \frac{1}{{160}}} \right)t = 1\]

OpenStudy (michele_laino):

please solve for t, keep in mind that: \[\Large \frac{1}{{240}} + \frac{1}{{160}} = \frac{{2 + 3}}{{480}} = \frac{5}{{480}}\]

OpenStudy (michele_laino):

after a substitution, we have: \[\Large \frac{5}{{480}}t = 1\] what is t?

OpenStudy (anonymous):

just simplify it?

OpenStudy (michele_laino):

hint: \[\Large t = \frac{{480}}{5} = ...?\]

OpenStudy (anonymous):

96

OpenStudy (anonymous):

I get that one

OpenStudy (anonymous):

So the last one

OpenStudy (michele_laino):

ok! question 1)

OpenStudy (michele_laino):

this question is similar to the question of candies, namely question 3). So the rate of working for the pump is W/40, when the pump works alone the working rate of the firehose is W/x, where x is the unknown quantity and W/60 is the working rate of the pump, when the pump and the firehose work together. So we can write: \[\Large \frac{W}{{40}} - \frac{W}{x} = \frac{W}{{60}}\] or: \[\Large \frac{1}{{40}} - \frac{1}{x} = \frac{1}{{60}}\]

OpenStudy (anonymous):

so the answer is 60?

OpenStudy (michele_laino):

no, since we have to compute the value of x

OpenStudy (michele_laino):

hint: \[\Large \frac{1}{x} = \frac{1}{{40}} - \frac{1}{{60}} = \frac{{3 - 2}}{{120}} = ...\]

OpenStudy (michele_laino):

please continue my computation

OpenStudy (anonymous):

1/120

OpenStudy (michele_laino):

ok! so we have: \[\Large x = \frac{{120}}{1} = ...{\text{minutes}}\]

OpenStudy (anonymous):

thanks very much I got 80%

OpenStudy (anonymous):

number 5 was wrong

OpenStudy (michele_laino):

are you sure?

OpenStudy (michele_laino):

I'm very sorry, the correct answer is option B. Sorry again

OpenStudy (anonymous):

That is totally fine

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

since we can write this: \[\Large \left( {\frac{W}{4} + \frac{W}{3}} \right)t = W\] and then: \[\Large \frac{1}{4} + \frac{1}{3} = \frac{1}{t}\] Sorry again

OpenStudy (michele_laino):

thanks for your comprehension!

OpenStudy (anonymous):

Thanks for helping me